Graph the given curve and find the area of a bounded region.


x= 3y^2 - 9

1
Expert's answer
2022-03-14T17:44:22-0400



"x=3y^2-9"

factoring;

"y=\sqrt{\frac{x+9}{3}}"

"y=(\frac{x+9}{3})^{\frac{1}{2}}"

to find the area we integrate the function with respect to x

"\int_{-9}^{\infin} (\frac{x+9}{3})^{\frac{1}{2}}dx"

the integral is divergent hence the area can not be found as it is unbounded.